A pharmacist needs to make 200 mL of a 10% hydrogen peroxide solution. She will make it from some 8% hydrogen peroxide solution and some 16% hydrogen peroxide solution. How much of the 8% solution and 16% solution should she use?
Found 2 solutions by josgarithmetic, stanbon:
Answer by josgarithmetic(39617)
Answer by stanbon(75887)

A pharmacist needs to make 200 mL of a 10% hydrogen peroxide solution. She will make it from some 8% hydrogen peroxide solution and some 16% hydrogen peroxide solution. How much of the 8% solution and 16% solution should she use?
-----------
Equation:
hyper + hyper = hyper
0.08x + 0.16(200-x) = 0.10*200
---------
8x + 16*200 - 16x = 10*200
---------
-8x = -6*200
-----
x = (3/4)200
x = 150 mL (amt. of 8% solution needed)
200-x = 50 mL (amt. of 16% solution needed)
---------------
Cheers,
Stan H.
------------